# Efficient designs for multivariate crossover trials

### v1

@@ Appendix B Efficiency under the generalized markov-type structure
$\begin{split}\bm{C}_{{d(s2)}(11)(1)}&=\begin{bmatrix}\bm{T}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\varOmega}_{1}\right)\bm{T}_{d}&-\bm{T}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\varOmega}_{2}\right)\bm{T}_{d}\\ -\bm{T}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\varOmega}_{2}\right)\bm{T}_{d}&\bm{T}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\varOmega}_{4}\right)\bm{T}_{d}\end{bmatrix},\\ \bm{C}_{{d(s2)}(12)(1)}&=\bm{C}^{{}^{\prime}}_{{d(s2)}(21)(1)}=\begin{bmatrix}\bm{T}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\varOmega}_{1}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}&-\bm{T}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\varOmega}_{2}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\\ -\bm{T}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\varOmega}_{2}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}&\bm{T}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\varOmega}_{4}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\end{bmatrix},\text{ and}\\ \bm{C}_{{d(s2)}(22)(1)}&=\begin{bmatrix}\bm{H}_{t}\bm{T}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\varOmega}_{1}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}&-\bm{H}_{t}\bm{T}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\varOmega}_{2}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\\ -\bm{H}_{t}\bm{T}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\varOmega}_{2}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}&\bm{H}_{t}\bm{T}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\varOmega}_{4}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\end{bmatrix}.\end{split}$
@@
$\begin{split}\bm{C}_{{d(s2)}(11)(1)}&=\begin{bmatrix}\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\varOmega}_{1}\right)\bm{T}_{d}&-\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\varOmega}_{2}\right)\bm{T}_{d}\\ -\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\varOmega}_{2}\right)\bm{T}_{d}&\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\varOmega}_{4}\right)\bm{T}_{d}\end{bmatrix},\\ \bm{C}_{{d(s2)}(12)(1)}&=\bm{C}^{{}^{\prime}}_{{d(s2)}(21)(1)}=\begin{bmatrix}\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\varOmega}_{1}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}&-\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\varOmega}_{2}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\\ -\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\varOmega}_{2}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}&\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\varOmega}_{4}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\end{bmatrix},\text{ and}\\ \bm{C}_{{d(s2)}(22)(1)}&=\begin{bmatrix}\bm{H}_{t}\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\varOmega}_{1}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}&-\bm{H}_{t}\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\varOmega}_{2}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\\ -\bm{H}_{t}\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\varOmega}_{2}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}&\bm{H}_{t}\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\varOmega}_{4}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\end{bmatrix}.\end{split}$ (S.96)
So, if $d\in\varOmega_{t,n,p}$ is a design uniform on periods, then from (S.95) and (S.96), we get that for $f_{1},f_{2}=1,2$, $\bm{{C}}_{d(s2)(f_{1}f_{2})(1)}=\bm{\tilde{C}}_{d(s2)(f_{1}f_{2})}$. Thus using the expression of the matrices $\bm{{C}}_{d(s2)}$ and $\bm{\tilde{C}}_{d(s2)}$ as given in Lemma S.3 and Lemma S.5, respectively, we get that $\bm{{C}}_{d(s2)}=\bm{\tilde{C}}_{d(s2)}$. ∎
@@
$\bm{V}_{R}^{-1}\bm{J}_{p\times p}\bm{V}_{R}^{-1}$. Then $tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{1}^{*}\bm{\psi}\right)>0$ and $tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)>0$. Here, $\bm{J}_{p\times p}=\bm{{1}}_{p}\bm{{1}}_{p}^{{}^{\prime}}$ and $\bm{H}_{p}=\bm{I}_{p}-\frac{1}{p}\bm{{1}}_{p}\bm{{1}}_{p}^{{}^{\prime}}$.
@@
$\bm{V}_{1}^{*}$ $=\bm{W}_{1}^{{}^{\prime}}\bm{W}_{1},$ (S.97)
where $\bm{W}_{1}=pr^{\perp}\left(\bm{V}_{1}^{-1/2}\bm{J}_{p\times p}\right)\bm{V}_{1}^{-1/2}$ and $\bm{J}_{p\times p}=\bm{{1}}_{p}\bm{{1}}_{p}^{{}^{\prime}}$. Since $\bm{H}_{p}=\bm{I}_{p}-\frac{1}{p}\bm{{1}}_{p}\bm{{1}}_{p}^{{}^{\prime}}$ is a symmetric and idempotent matrix, $tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{1}^{*}\bm{\psi}\right)=tr\left(\bm{W}_{2}^{{}^{\prime}}\bm{W}_{2}\right)\geq 0$, where $\bm{W}_{2}=\bm{W}_{1}\bm{\psi}\bm{H}_{p}$. It can be easily seen that $tr\left(\bm{W}_{2}^{{}^{\prime}}\bm{W}_{2}\right)=\bm{{0}}_{p\times p}$ if and only if $\bm{W}_{2}=\bm{{0}}_{p\times p}$. Now, suppose $tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{1}^{*}\bm{\psi}\right)=0$. So, $\bm{W}_{1}\bm{\psi}\bm{H}_{p}=\bm{{0}}_{p\times p}$, and thus from (S.97), we get
@@
where $\left(\bm{V}_{1}^{*}\right)_{l_{1},l_{2}}$ denotes the $(l_{1},l_{2})^{th}$ element of $\bm{V}_{1}^{*}$. Using the property that the row sums of $\bm{V}_{1}^{*}$ is 00 in (S.99), we get that $\bm{V}_{1}^{*}=\bm{{0}}_{p\times p}$. Thus by calculation we get
@@
The above equation (S.100) leads to a contradiction. Hence, our assumption that
$tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{1}^{*}\bm{\psi}\right)=0$ is wrong. Thus we get, $tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{1}^{*}\bm{\psi}\right)>0$.

### v2

@@ Appendix B Efficiency under the generalized Markov-type structure
$\begin{split}\bm{C}_{{d(s2)}(11)(1)}&=\begin{bmatrix}\bm{T}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\mathit{\Omega}}_{1}\right)\bm{T}_{d}&-\bm{T}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\mathit{\Omega}}_{2}\right)\bm{T}_{d}\\ -\bm{T}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\mathit{\Omega}}_{2}\right)\bm{T}_{d}&\bm{T}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\mathit{\Omega}}_{4}\right)\bm{T}_{d}\end{bmatrix},\\ \bm{C}_{{d(s2)}(12)(1)}&=\bm{C}^{{}^{\prime}}_{{d(s2)}(21)(1)}=\begin{bmatrix}\bm{T}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\mathit{\Omega}}_{1}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}&-\bm{T}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\mathit{\Omega}}_{2}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\\ -\bm{T}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\mathit{\Omega}}_{2}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}&\bm{T}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\mathit{\Omega}}_{4}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\end{bmatrix},\text{ and}\\ \bm{C}_{{d(s2)}(22)(1)}&=\begin{bmatrix}\bm{H}_{t}\bm{T}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\mathit{\Omega}}_{1}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}&-\bm{H}_{t}\bm{T}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\mathit{\Omega}}_{2}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\\ -\bm{H}_{t}\bm{T}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\mathit{\Omega}}_{2}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}&\bm{H}_{t}\bm{T}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\mathit{\Omega}}_{4}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\end{bmatrix}.\end{split}$
@@
$\begin{split}\bm{C}_{{d(s2)}(11)(1)}&=\begin{bmatrix}\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\mathit{\Omega}}_{1}\right)\bm{T}_{d}&-\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\mathit{\Omega}}_{2}\right)\bm{T}_{d}\\ -\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\mathit{\Omega}}_{2}\right)\bm{T}_{d}&\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\mathit{\Omega}}_{4}\right)\bm{T}_{d}\end{bmatrix},\\ \bm{C}_{{d(s2)}(12)(1)}&=\bm{C}^{{}^{\prime}}_{{d(s2)}(21)(1)}=\begin{bmatrix}\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\mathit{\Omega}}_{1}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}&-\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\mathit{\Omega}}_{2}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\\ -\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\mathit{\Omega}}_{2}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}&\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\mathit{\Omega}}_{4}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\end{bmatrix},\text{ and}\\ \bm{C}_{{d(s2)}(22)(1)}&=\begin{bmatrix}\bm{H}_{t}\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\mathit{\Omega}}_{1}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}&-\bm{H}_{t}\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\mathit{\Omega}}_{2}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\\ -\bm{H}_{t}\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\mathit{\Omega}}_{2}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}&\bm{H}_{t}\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\mathit{\Omega}}_{4}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\end{bmatrix}.\end{split}$ (S.49)
So, if $d\in\Omega_{t,n,p}$ is a design uniform on periods, then from (S.48) and (S.49), we get that for $f_{1},f_{2}=1,2$, $\bm{{C}}_{d(s2)(f_{1}f_{2})(1)}=\bm{\tilde{C}}_{d(s2)(f_{1}f_{2})}$. Thus using the expression of the matrices $\bm{{C}}_{d(s2)}$ and $\bm{\tilde{C}}_{d(s2)}$ as given in Lemma S.3 and Lemma S.5, respectively, we get that $\bm{{C}}_{d(s2)}=\bm{\tilde{C}}_{d(s2)}$. ∎
@@
$\bm{V}_{R}^{-1}\bm{J}_{p\times p}\bm{V}_{R}^{-1}$. Then $tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{1}^{*}\bm{\psi}\right)>0$ and $tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)>0$.
@@
$\bm{V}_{1}^{*}$ $=\bm{L}_{1}^{{}^{\prime}}\bm{L}_{1},$ (S.50)
where $\bm{L}_{1}=pr^{\perp}\left(\bm{V}_{1}^{-1/2}\bm{J}_{p\times p}\right)\bm{V}_{1}^{-1/2}$. Since $\bm{H}_{p}$ is a symmetric and idempotent matrix,
$tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{1}^{*}\bm{\psi}\right)=tr\left(\bm{L}_{2}^{{}^{\prime}}\bm{L}_{2}\right)\geq 0$, where $\bm{L}_{2}=\bm{L}_{1}\bm{\psi}\bm{H}_{p}$. It can be easily seen that $tr\left(\bm{L}_{2}^{{}^{\prime}}\bm{L}_{2}\right)=\bm{{0}}_{p\times p}$ if and only if $\bm{L}_{2}=\bm{{0}}_{p\times p}$. Now, suppose $tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{1}^{*}\bm{\psi}\right)=0$. So, $\bm{L}_{1}\bm{\psi}\bm{H}_{p}=\bm{{0}}_{p\times p}$ and thus from (S.50), we get
@@
where $\left(\bm{V}_{1}^{*}\right)_{l_{1},l_{2}}$ denotes the $(l_{1},l_{2})^{th}$ element of $\bm{V}_{1}^{*}$. Using the property that the row sums of $\bm{V}_{1}^{*}$ is 00 in (S.52), we get that $\bm{V}_{1}^{*}=\bm{{0}}_{p\times p}$. Thus, by calculation we get
@@
The above equation (S.53) leads to a contradiction. Hence our assumption that
$tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{1}^{*}\bm{\psi}\right)=0$ is wrong. Thus, $tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{1}^{*}\bm{\psi}\right)>0$.